News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 28 September 2001 12.00pm EDT

During last week the observing schedule was interrupted after the Observatory entered a bright star hold at 7.48am on Sept 22. The entry followed a 2.5 hour observation of the dark moon performed on gyro-hold. Analysis of the dump data indicates that a gyro bias error of ~170 arcsec accumulated during the gyro hold and subsequent maneuver. The error exceeded the size of the aspect camera search box during the next star acquisition and resulted in the bright star hold. Data indicates that the gyro performance remains consistent with past performance, however a detailed analysis will be performed by the gyro vendor. Fine attitude was recovered at 11pm EDT on Sept 22, and loads were uplinked at 2pm on Sept 23. The observations of Comet 19P/Borrelly, Puppis N Knot and 1RXS J1708-4009 were impacted by the bright star hold.

The loads were interrupted a second time at 10.29am on Sept 24 by a significant solar radiation event that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and assessment of the dump data confirmed that the trigger was due to high proton counts in the EPHIN E1300 channel exceeding the pre-defined threshold of 10 counts/cm^2/s/sr. The solar levels have remained high throughout the week but are projected to reach an acceptable level during Sept 29. The loads are planned to resume at 12.50am on Sept 30. The 10 observing targets impacted by the radiation event will be rescheduled in a future load.

The schedule for next week is shown below.

------------------------------------------
SGR 1627-41             ACIS-S       Sep 30
IRAS 23128-5919         ACIS-S
Puppis A shock front    ACIS-S       Oct 1
Radiation Belts
CAL 83                  HRC-S/LETG   Oct 2
UV CETI,GLIESE 65       HRC-S/LETG
IRAS 05189-2524         ACIS-S       Oct 3
Puppis bright E knot    ACIS-S
Radiation Belts                      Oct 4
ABELL 496               ACIS-S
M31 Nucleus             ACIS-S
M31 Nucleus             ACIS-I       Oct 5
A2556                   ACIS-S
SAX J1753.5-2349        ACIS-S
XTE1650-500             ACIS-S/HETG
NGC6093                 ACIS-S       Oct 6
Radiation Belts
RCW103                  ACIS-S       Oct 7
PKS 2135-14             ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports